# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 4.1.3 Simulation Study: The Brock–Mirman Economy
The VI iteration count follows from the contraction bound. Each iteration reduces the Bellman residual $\|V_{k+1}-V_{k}\|_{\infty}$ by the factor $\beta=0.96$: a 4% reduction per step. With initial residual $\|TV_{0}-V_{0}\|_{\infty}$ determined by the model primitives and tolerance $\epsilon=10^{-10}$, the required iteration count is $k=\lceil\log(\epsilon/\|TV_{0}-V_{0}\|_{\infty})/\log\beta\rceil=567$. At $\beta=0.99$, the weaker contraction yields approximately four times as many iterations for the same tolerance, because $\log(0.96)/\log(0.99)\approx 4$.
Figure 2(a)–(b) provide a geometric interpretation for a scalar Bellman equation with three policies. Each policy operator $T^{\pi_{i}}V=r^{\pi_{i}}+\gamma_{i}V$ is affine; the Bellman operator $T=\max_{i}T^{\pi_{i}}$ is their upper envelope, a convex piecewise-linear function. Panel (a) shows VI: the staircase iterates $V_{k+1}=TV_{k}$ by alternating between the $T$ curve and the $45^{\circ}$ line, converging at the linear rate $\gamma$. Panel (b) shows PI: at each iterate, the algorithm identifies the active policy operator and solves for its fixed point on the $45^{\circ}$ line, jumping directly to the intersection. This is a Newton step: each $T^{\pi_{k}}$ is a supporting hyperplane to $T$ at $V_{k}$, and the fixed point of the linearization is the Newton iterate. The scalar picture extends to $\mathbb{R}^{n}$: the affine operator $T^{\pi_{k}}V=r^{\pi_{k}}+\gamma P^{\pi_{k}}V$ supports $T$ at $V_{k}$, and its fixed point $(I-\gamma P^{\pi_{k}})^{-1}r^{\pi_{k}}$ is the Newton iterate from equation (28). Finite termination follows because $T$ has finitely many affine pieces; the iteration count depends on the number of policy switches, not the state-space dimension.
[…]
Figure 2: The Brock–Mirman economy ($\alpha=0.36$, $\beta=0.96$, 1,000 states). (a) Value iteration on a scalar Bellman equation: the staircase iterates $V_{k+1}=TV_{k}$, converging at the linear rate $\gamma$. (b) Policy iteration as Newton’s method: each step solves for the fixed point of the active policy operator $T^{\pi_{k}}$, jumping to the tangent line’s intersection with the diagonal. (c) Sup-norm error $\|V_{k}-V^{*}\|_{\infty}$ for the discretized model; VI requires 567 iterations, PI converges in $1$.
@@ 4.2.2 Q-Learning and SARSA
The Q-learning update, upon observing transition $(s_{t},a_{t},r_{t},s_{t+1})$, is:
@@
Convergence requires two conditions. Exploration (visiting all state-action pairs infinitely often) ensures identification; every state-action pair must be sampled to learn its value. The Robbins-Monro step-size conditions ($\sum_{t}\alpha_{t}=\infty$, $\sum_{t}\alpha_{t}^{2}<\infty$) balance tracking versus noise suppression. Watkins and Dayan (1992) and Jaakkola et al. (1994) formalize these; Tsitsiklis (1994) provides the general framework.
@@ 4.2.3 Rollout, Lookahead, and AlphaZero
This is one step of policy iteration starting from $\mu$: the policy improvement theorem guarantees $V^{\mu_{R}}(s)\geq V^{\mu}(s)$ for all $s$, with strict inequality unless $\mu$ is already optimal (Bertsekas, 2021). [^50] Bertsekas uses cost-minimization notation throughout his work, writing $\min$ where standard RL uses $\max$ and defining value as accumulated cost rather than reward. I translate to the reward-maximization convention used elsewhere in this chapter; the mathematics are equivalent with reversed inequalities. Given an arbitrary approximate value function $\tilde{V}$ (not necessarily the value of any policy), the one-step lookahead policy selects
@@ 4.3 The Central Challenge: The Deadly Triad
The preceding results assume tabular representations, namely, a separate value for each state-action pair. In practice, state spaces are too large for lookup tables (Go has $10^{170}$ states; most economic models have continuous state variables). Practitioners must therefore combine three ingredients: function approximation (to handle large state spaces), bootstrapping (to learn from single transitions rather than complete episodes), and off-policy learning (to learn about the optimal policy while exploring, or to reuse old data). Each is desirable in isolation. Their interaction, known as the deadly triad (Sutton and Barto, 2018, Ch. 11), is the central open problem in reinforcement learning theory.
@@ 4.3.1 The Projected Bellman Operator
With function approximation $V(s)\approx\phi(s)^{\top}\theta$, the parameter vector $\theta$ is shared across states. Updating $\theta$ to improve the value estimate at one state simultaneously changes the estimate at every other state. The algorithm can no longer apply the Bellman operator $T^{\pi}$ to each state independently; instead, it applies $T^{\pi}$ to compute a target, then projects the result back onto the function space (the span of the features $\phi$). The composed operator is $\Pi T^{\pi}$, where $\Pi$ denotes this projection (Tsitsiklis and Van Roy, 1997). Convergence of the approximate iteration $\theta_{k+1}=\Pi T^{\pi}\theta_{k}$ requires $\Pi T^{\pi}$ to be a contraction.
In the on-policy setting, $\Pi$ minimizes squared error weighted by $d^{\pi}$, the stationary distribution of the policy being evaluated: $\Pi V=\arg\min_{\hat{V}\in\text{span}(\Phi)}\|V-\hat{V}\|_{d^{\pi}}$, where $\|V\|_{d^{\pi}}^{2}=\sum_{s}d^{\pi}(s)V(s)^{2}$. The Bellman operator $T^{\pi}$ is a $\gamma$-contraction in the same $d^{\pi}$-norm. Because both operators use the same norm, the projection $\Pi$ is orthogonal: the residual $V-\Pi V$ is perpendicular to the approximation subspace. The Pythagorean theorem then gives $\|\Pi V\|_{d^{\pi}}^{2}+\|V-\Pi V\|_{d^{\pi}}^{2}=\|V\|_{d^{\pi}}^{2}$, so $\|\Pi V\|_{d^{\pi}}\leq\|V\|_{d^{\pi}}$. [^55] The same argument in $\mathbb{R}^{n}$: projecting a vector onto a subspace never makes it longer. This is the geometric content of the Cauchy-Schwarz inequality. In the function-approximation setting, the “vector” is a value function, the “subspace” is the span of features, and “length” is the $d^{\pi}$-weighted $L^{2}$ norm. The projection cannot expand distances. The composition therefore contracts:

### 15 Mar 2026 v2

@@ 5.1.3 Simulation Study: The Brock–Mirman Economy
The VI iteration count follows from the contraction bound: each iteration reduces the Bellman residual by the factor $\beta=0.96$, requiring $k=\lceil\log(\epsilon/\|TV_{0}-V_{0}\|_{\infty})/\log\beta\rceil=567$ iterations for tolerance $\epsilon=10^{-10}$. [^50] At $\beta=0.99$, the weaker contraction yields approximately four times as many iterations, since $\log(0.96)/\log(0.99)\approx 4$.
Figure 3(a)–(b) provide a geometric interpretation for a scalar Bellman equation with three policies. Each policy operator $T^{\pi_{i}}V=r^{\pi_{i}}+\gamma_{i}V$ is affine; the Bellman operator $T=\max_{i}T^{\pi_{i}}$ is their upper envelope, a convex piecewise-linear function. Panel (a) shows VI: the staircase iterates $V_{k+1}=TV_{k}$ by alternating between the $T$ curve and the $45^{\circ}$ line, converging at the linear rate $\gamma$. Panel (b) shows PI: at each iterate, the algorithm identifies the active policy operator and solves for its fixed point on the $45^{\circ}$ line, jumping directly to the intersection. This is a Newton step, where each $T^{\pi_{k}}$ is a supporting hyperplane to $T$ at $V_{k}$, and the fixed point of the linearization is the Newton iterate. The scalar picture extends to $\mathbb{R}^{n}$: the affine operator $T^{\pi_{k}}V=r^{\pi_{k}}+\gamma P^{\pi_{k}}V$ supports $T$ at $V_{k}$, and its fixed point $(I-\gamma P^{\pi_{k}})^{-1}r^{\pi_{k}}$ is the Newton iterate from equation (33). Finite termination follows because $T$ has finitely many affine pieces; the iteration count depends on the number of policy switches, not the state-space dimension.
Table 2 and Figure 3 confirm the theory. VI requires 567 iterations at rate $\beta^{n}=0.96^{n}$; PI converges in 11, a $50\times$ reduction predicted by the Newton interpretation. The Manne (1960) LP recovers the same value function to machine precision ($\|V_{\text{LP}}-V_{\text{VI}}\|_{\infty}<10^{-8}$). [^51] PI wall-clock time scales favorably: at $n_{k}=200$ grid points, PI finishes in 0.057 seconds versus VI’s 2.4 seconds, because the $O(n^{3})$ per-iteration cost of policy evaluation is offset by 7–10 total iterations versus 567.
Figure 3: The Brock–Mirman economy ($\alpha=0.36$, $\beta=0.96$, 1,000 states). (a) Value iteration on a scalar Bellman equation. The staircase iterates $V_{k+1}=TV_{k}$, converging at the linear rate $\gamma$. (b) Policy iteration as Newton’s method. Each step solves for the fixed point of the active policy operator $T^{\pi_{k}}$, jumping to the tangent line’s intersection with the diagonal. (c) Sup-norm error $\|V_{k}-V^{*}\|_{\infty}$ for the discretized model; VI requires 567 iterations, PI converges in $1$.
@@ 5.2.2 Q-Learning and SARSA
The Q-learning update, upon observing transition $(s_{t},a_{t},r_{t},s_{t+1})$, is
@@
Convergence requires two conditions. Exploration (visiting all state-action pairs infinitely often) ensures identification. The Robbins-Monro step-size conditions ($\sum_{t}\alpha_{t}=\infty$, $\sum_{t}\alpha_{t}^{2}<\infty$) balance tracking versus noise suppression. Watkins and Dayan (1992) and Jaakkola et al. (1994) formalize these; Tsitsiklis (1994) provides the general framework.
@@
5.2.3 Multi-Step Returns and TD($\lambda$)
The updates in (36) bootstrap from a single successor state. More generally, one can bootstrap from $n$ steps ahead. The $n$-step return is
$G_{t}^{(n)}=\sum_{k=0}^{n-1}\gamma^{k}R_{t+k+1}+\gamma^{n}V(S_{t+n}).$ (38)
Setting $n=1$ recovers the TD(0) target; letting $n\to\infty$ (or reaching a terminal state) gives the Monte Carlo return.
The $\lambda$-return (Sutton, 1988) averages all $n$-step returns with geometrically decaying weights:
$G_{t}^{\lambda}=(1-\lambda)\sum_{n=1}^{\infty}\lambda^{n-1}G_{t}^{(n)},\qquad\lambda\in[0,1].$ (39)
This is the forward view: at time $t$, look forward at all possible truncation horizons and take a weighted average. The parameter $\lambda$ controls a bias-variance tradeoff: lower $\lambda$ gives higher bias (more bootstrapping) but lower variance; higher $\lambda$ gives lower bias but higher variance, since more of the update relies on stochastic returns rather than value estimates.
The forward view requires waiting until the end of the episode to compute $G_{t}^{\lambda}$. The backward view computes the same total update incrementally. At each step, compute one TD error $\delta_{t}=R_{t+1}+\gamma V(S_{t+1})-V(S_{t})$ and distribute it to all states via an eligibility trace:
$e_{t}(s)=\gamma\lambda\,e_{t-1}(s)+\mathbbm{1}\{s=S_{t}\},\qquad V(s)\leftarrow V(s)+\alpha\,\delta_{t}\,e_{t}(s).$ (40)
The trace $e_{t}(s)$ acts as a fading memory of recently visited states: it spikes when $s$ is visited and decays by $\gamma\lambda$ per step. Each TD error $\delta_{t}$ updates every state in proportion to its current trace, enabling $O(|\mathcal{S}|)$ per-step credit assignment without storing trajectories. [^59] The forward and backward views produce identical total weight changes over a complete episode (Sutton, 1988). The backward view is preferred in practice because it operates online (updating after each transition) rather than requiring the full episode to be stored. Practical variants include replacing traces ($e_{t}(s)=1$ when $s=S_{t}$, capping the trace at 1 instead of accumulating), Dutch traces (van Seijen et al., 2016), and off-policy extensions such as Retrace($\lambda$) (Munos et al., 2016) and V-trace (Espeholt et al., 2018). The historical development of eligibility traces is discussed in Section 3.
Under linear function approximation, TD($\lambda$) converges to a unique fixed point with approximation error bounded by $\frac{1-\lambda\gamma}{\sqrt{1-\gamma^{2}}}$ times the best-in-class error (Equation 46 and the bound in Section 5.3; Tsitsiklis and Van Roy, 1997). Higher $\lambda$ tightens this bound, approaching the projection of $V^{\pi}$ as $\lambda\to 1$.
5.2.4 Simulation Study: Credit Assignment in a Corridor
A 20-state deterministic corridor ($s\in\{0,\ldots,19\}$, action: move right, reward +1+1 only at the terminal state $s=19$, $\gamma=0.99$) isolates the credit-assignment mechanism. The true value function is $V^{*}(s)=\gamma^{19-s}$. TD($\lambda$) performs policy evaluation for four values of $\lambda$ across 20 seeds and 200 episodes.
Table 3 and Figure 4 show that higher $\lambda$ propagates the sparse terminal reward signal backward through the corridor faster: TD($\lambda=1$) reaches RMSVE <0.05<0.05 in fewer episodes than TD(0), which must wait for many episodes before the reward signal diffuses back to early states through one-step bootstrapping alone.
Figure 4: RMSVE vs. episodes for TD($\lambda$) on the 20-state corridor. Shaded regions show $\pm 1$ SE over 20 seeds. Higher $\lambda$ propagates the terminal reward faster.
Table 3: TD($\lambda$) on 20-state corridor. Mean $\pm$ SE over 20 seeds, 200 episodes, $\gamma=0.99$, $\alpha=0.05$.
$\lambda$ Final RMSVE Episodes to RMSVE <0.05<0.05
0.0 0.3944 $\pm$ 0.0056 >200>200
0.4 0.1843 $\pm$ 0.0028 >200>200
0.8 0.0086 $\pm$ 0.0001 136 $\pm$ 1
1.0 0.0091 $\pm$ 0.0000 48 $\pm$ 0
5.2.5 Finite-Sample Theory of Fitted Methods
Fitted Q-Iteration and Fitted Value Iteration (Definition 4.1.10, Section 4.1.10) replace exact Bellman applications with projected regression steps. The projection introduces approximation error that compounds across iterations.
Define the inherent Bellman approximation error
$\varepsilon_{\mathrm{approx}}=\inf_{f\in\mathcal{F}}\|\mathcal{T}f-f\|_{p,\mu},$ (41)
the smallest residual achievable when the Bellman operator maps any element of $\mathcal{F}$ back to itself. Munos and Szepesvári (2008) show that after $K$ iterations with $N$ i.i.d. samples per iteration,
$\|V_{K}-V^{*}\|_{p,\rho}\;\leq\;C_{\rho,\mu}\left[\gamma^{K}\|V_{0}-V^{*}\|_{p,\rho}+\frac{\varepsilon_{\mathrm{approx}}}{(1-\gamma)^{2}}+O\!\left(\frac{1}{\sqrt{N}}\right)\right],$ (42)
where $C_{\rho,\mu}$ is a concentrability coefficient bounding the ratio of future-state distributions under the evaluation distribution $\rho$ relative to the data distribution $\mu$. [^60] The concentrability coefficient $C_{\rho,\mu}$ measures how well the data distribution $\mu$ covers future states reachable under optimal policies from $\rho$. When $\mu$ is the state-action distribution of the optimal policy itself, $C_{\rho,\mu}=1$. Distribution mismatch, common when batch data comes from a sub-optimal behavior policy, inflates $C_{\rho,\mu}$ and worsens the bound. Antos et al. (2008) extend these results to continuous action spaces and single-trajectory data. Three terms drive the error: the geometric decay $\gamma^{K}$ (initialization bias), the approximation error $(1-\gamma)^{-2}\varepsilon_{\mathrm{approx}}$ (bias from function class), and the estimation error $O(1/\sqrt{N})$ (variance from finite samples). When $\mathcal{F}$ contains $V^{*}$ exactly, $\varepsilon_{\mathrm{approx}}=0$ and the bound recovers exact convergence as $K\to\infty$. The $(1-\gamma)^{-2}$ amplification, one factor of $(1-\gamma)^{-1}$ more than in tabular Q-learning, reflects error accumulation across approximate DP steps: each regression step introduces bias, and this bias compounds over $K$ iterations.
Both algorithms solve projected Bellman equations. When $V^{*}\in\mathrm{span}(\Phi)$ exactly, FVI converges to $V^{*}$ in a single projected iteration, since the normal equations (23) recover $\theta_{V}^{*}$ satisfying $\Phi\theta_{V}^{*}=V^{*}$. For FQI, the per-action Q-functions satisfy $Q^{*}(s,a^{*}(s))=V^{*}(s)$ at the optimal action $a^{*}(s)$, so when $Q^{*}(\cdot,a)$ is also representable in $\mathrm{span}(\Phi)$ for each $a$, FQI recovers consistent value estimates $\phi(s)^{\top}\theta_{a^{*}(s)}^{*}=\phi(s)^{\top}\theta_{V}^{*}$ at convergence. Whether FQI succeeds therefore depends on the geometry of the problem: when $Q^{*}(\cdot,a)\notin\mathrm{span}(\Phi)$, as on the Brock–Mirman economy, where the per-action Q-function requires fractional-power terms $k^{-n\alpha}$ outside the log-polynomial span, FQI stalls at error 1.65 while FVI converges to 0.001; when $Q^{*}(x,u)$ is exactly quadratic in $(x,u)$, as on the linear-quadratic control problem (Section 5.2.6), both FVI and FQI converge to near-zero error. Section 5.2.7 shows that replacing the linear basis with a nonlinear parametric model that matches the log-Cobb–Douglas structure of $Q^{*}$ restores FQI convergence on the Brock–Mirman economy.
5.2.6 Simulation Study: Fitted Methods on Linear-Quadratic Control
Linear-quadratic control (LQC) is a setting where both $V^{*}$ and $Q^{*}$ are quadratic polynomials, so the fitted method comparison is analytically transparent. The model has scalar state $x\in[-4,4]$ and action $u\in[-2,2]$, deterministic dynamics $x^{\prime}=ax+bu$ with $a=0.5$, $b=1.0$, reward $r(x,u)=-(x^{2}+u^{2})$, and discount $\gamma=0.95$. The parameters are chosen so that $x^{\prime}=0.5x+u\in[-4,4]$ whenever $(x,u)\in[-4,4]\times[-2,2]$, making the grid strictly invariant. The optimal value function satisfies $V^{*}(x)=-Px^{2}$, where $P$ solves $\gamma b^{2}P^{2}+P(1-\gamma(a^{2}+b^{2}))-1=0$, yielding $P\approx 1.129$. The optimal Q-function is $Q^{*}(x,u)=-(1+\gamma Pa^{2})x^{2}-2\gamma Pab\,xu-(1+\gamma Pb^{2})u^{2}\approx-1.268x^{2}-1.073xu-2.073u^{2}$, which lies exactly in $\mathrm{span}\{x^{2},xu,u^{2}\}\subset\mathrm{span}\{x,x^{2},u,u^{2},xu\}$. FVI uses state features $\phi_{V}(x)=[x,x^{2}]^{\top}\in\mathbb{R}^{2}$ (no intercept, since $V^{*}(0)=0$); FQI uses state-action features $\phi_{Q}(x,u)=[x,x^{2},u,u^{2},xu]^{\top}\in\mathbb{R}^{5}$. Both use a 301-point state grid and 201-point action grid. DQN uses a two-layer network of 64 units per layer with ReLU activations, an experience replay buffer of 50,000 transitions, a hard target-network update every 500 gradient steps, and rewards scaled by a factor of $1/20$ to stabilize training.
Both methods recover the analytical solution to machine precision (Table 4, Figure 5): FVI and FQI converge in under 10 iterations with errors below $10^{-3}$, matching the known polynomial structure of $Q^{*}$. DQN also converges (error $5.6\times 10^{-1}$ after 100,000 gradient steps) with no prior knowledge of the feature basis. The contrast with Brock–Mirman is exact: FQI succeeds here because $Q^{*}(x,u)\in\mathrm{span}(\Phi_{Q})$, while it fails on Brock–Mirman because $Q^{*}(\cdot,a)\notin\mathrm{span}(\Phi)$.
Method Iterations Error vs $V^{*}$ $P$ (recovered) Key coefficient
Exact VI (discrete) 25 1.12e-03
FVI 9 3.23e-04 1.1294 $\hat{\theta}_{V}^{x^{2}}=-1.1294$
FQI 10 9.37e-05 1.2682 $\hat{\theta}_{Q}^{xu}=-1.0729$
DQN ($2\times 64$ ReLU) 100000 5.64e-01
Analytical ($V^{*}=-Px^{2}$) 0 1.1294 $c_{xu}=-1.0729$
Table 4: Fitted weights and convergence metrics for FVI, FQI, and DQN on linear-quadratic control. The FVI $x^{2}$ coefficient recovers the Riccati solution $P\approx 1.129$. The FQI quadratic coefficients match the analytical $Q^{*}$ to four decimal places. FVI and FQI achieve max error below $10^{-3}$ against the analytical $V^{*}$; DQN achieves error $5.6\times 10^{-1}$ after 100,000 gradient steps with no feature basis specified.
Figure 5: LQC convergence of FVI and FQI (left), DQN learning curve (middle), and value function recovery for all three methods (right). FVI and FQI reduce $\|V_{k}-V^{*}\|_{\infty}$ to near-zero in under 10 iterations, exploiting the known polynomial structure of $Q^{*}$. DQN declines from error 6.7 to 0.56 over 100,000 gradient steps with no feature basis specified.
5.2.7 Simulation Study: Basis Representability on the Brock–Mirman Economy
[…]
Table 5 and Figure 6 confirm the diagnosis: basis representability, not algorithmic failure. FVI converges near the projection floor of the log-polynomial basis; linear FQI stalls at error 1.65, confirming $Q^{*}(\cdot,a)\notin\mathrm{span}(\Phi)$. Oracle-FQI and NLLS-FQI, using the structurally correct log-consumption feature, match exact VI with error below $10^{-4}$. NLLS-FQI recovers $\hat{\alpha}=0.3600$ in a single iteration, demonstrating that the same FQI algorithm succeeds when the function class contains $Q^{*}$.
Method $\|V-V^{*}\|_{\infty}$ $\|V-V^{*}\|_{\text{rms}}$ Policy agreement (%) Iterations
Exact VI 0.0000 0.0000 98.0
FVI (linear) 0.0010 0.0008 99.0 341
FQI (linear) 1.6521 1.4805 4.0 339
Oracle-FQI 0.0000 0.0000 98.0 341
NLLS-FQI ($\hat{\alpha}=0.3600$) 0.0000 0.0000 98.0 341
Table 5: Convergence metrics for five methods on the Brock–Mirman economy ($N_{K}=50$, $N_{Z}=2$, $\gamma=0.96$). Policy agreement is measured against the closed-form optimal policy $k^{\prime}=\alpha\beta zk^{\alpha}$.
Figure 6: Left: convergence of $\|V_{k}-V^{*}\|_{\infty}$ for FVI, linear FQI, Oracle-FQI, and NLLS-FQI on the Brock–Mirman economy. Right: NLLS-FQI estimated $\alpha$ trajectory, converging from $\alpha_{0}=0.5$ to the true $\alpha=0.36$ in one iteration.
@@ 5.2.8 Rollout, Lookahead, and AlphaZero
This is one step of policy iteration starting from $\mu$: the Policy Improvement Theorem (Theorem 1) guarantees $V^{\mu_{R}}(s)\geq V^{\mu}(s)$ for all $s$, with strict inequality unless $\mu$ is already optimal (Bertsekas, 2021). [^62] Bertsekas uses cost-minimization notation throughout his work, writing $\min$ where standard RL uses $\max$ and defining value as accumulated cost rather than reward. I translate to the reward-maximization convention used elsewhere in this chapter; the mathematics are equivalent with reversed inequalities. Given an arbitrary approximate value function $\tilde{V}$ (not necessarily the value of any policy), the one-step lookahead policy selects
@@ 5.3 The Central Challenge: The Deadly Triad
State spaces are too large for lookup tables (Go has $10^{170}$ states; most economic models have continuous state variables). Practitioners must combine three ingredients: function approximation (to handle large state spaces), bootstrapping (to learn from single transitions rather than complete episodes), and off-policy learning (to learn about the optimal policy while exploring, or to reuse old data). Each is desirable in isolation. Their interaction, known as the deadly triad (Sutton and Barto, 2018, Ch. 11), is the central open problem in reinforcement learning theory.
Off-policy learning is preferred for three reasons. First, sample efficiency: transitions collected under any behavioral policy can be reused to evaluate or improve a different target policy, amortizing the cost of data collection. Discarding data because it was generated by a superseded policy is wasteful. Second, exploration and exploitation separate cleanly: the agent can follow an exploratory policy (e.g., $\varepsilon$-greedy) to ensure adequate state-space coverage while simultaneously learning the optimal deterministic policy. On-policy methods such as SARSA entangle the two, learning the value of the exploratory policy rather than the optimal one. Third, off-policy evaluation answers “what would have happened under policy $\pi$?” from data generated by policy $\mu$, the counterfactual question at the heart of policy comparison.
@@ 5.3.1 The Projected Bellman Operator
With function approximation $V(s)\approx\phi(s)^{\top}\theta$, the parameter vector $\theta$ is shared across states. Updating $\theta$ to improve the value estimate at one state simultaneously changes the estimate at every other state. The algorithm can no longer apply the Bellman operator $T^{\pi}$ to each state independently; instead, it applies $T^{\pi}$ to compute a target, then projects the result back onto the function space (the span of the features $\phi$). The composed operator is $\Pi T^{\pi}$, the projected Bellman operator, where $\Pi$ denotes this projection (Tsitsiklis and Van Roy, 1997). Convergence of the approximate iteration $\theta_{k+1}=\Pi T^{\pi}\theta_{k}$ requires $\Pi T^{\pi}$ to be a contraction.
In the on-policy setting, $\Pi$ minimizes squared error weighted by $d^{\pi}$, the stationary distribution of the policy being evaluated, so $\Pi V=\arg\min_{\hat{V}\in\text{span}(\Phi)}\|V-\hat{V}\|_{d^{\pi}}$, where $\|V\|_{d^{\pi}}^{2}=\sum_{s}d^{\pi}(s)V(s)^{2}$. The Bellman operator $T^{\pi}$ is a $\gamma$-contraction in the same $d^{\pi}$-norm. Because both operators use the same norm, the projection $\Pi$ is orthogonal, meaning the residual $V-\Pi V$ is perpendicular to the approximation subspace. The Pythagorean theorem then gives $\|\Pi V\|_{d^{\pi}}^{2}+\|V-\Pi V\|_{d^{\pi}}^{2}=\|V\|_{d^{\pi}}^{2}$, so $\|\Pi V\|_{d^{\pi}}\leq\|V\|_{d^{\pi}}$. [^67] The same argument holds in $\mathbb{R}^{n}$. Projecting a vector onto a subspace never makes it longer. This is the geometric content of the Cauchy-Schwarz inequality. In the function-approximation setting, the “vector” is a value function, the “subspace” is the span of features, and “length” is the $d^{\pi}$-weighted $L^{2}$ norm. The projection cannot expand distances. The composition therefore contracts: